# Guiding an Automatic Speech Recognition Decoder using Large Language Models

### v1

@@ Abstract
Automatic Speech Recognition (ASR) consists of an acoustic model (AM) and a language model (LM). The AM estimates the probability of an acoustic signal based on a sequence of linguistic units, typically phones, characters, or tokens, while the LM assesses the likelihood of a specific sequence of words or tokens. Although Large Language Models (LLMs) have demonstrated significant potential across various tasks, integrating them into ASR remains an open challenge. By decomposing the maximum a posteriori (MAP) estimator of words (or tokens) given the acoustic signal, we derive an iterative procedure that facilitates a novel integration of the AM and LLM, while maintaining their separability. This approach enables each component to be independently trained and improved using its own data, thereby maximizing the system’s performance by leveraging the strengths of both models without requiring joint optimization. We illustrate the effectiveness of our method in comparison to three language models: N-gram, GCNN, and TransformerLM across multiple datasets spanning various speech styles, including ALLSSTAR, WSJ0, and TED-LIUM 3. Our experiments involved two acoustic models (wav2vec 2.0 and HuBERT) and three LLMs (GPT-2, LLaMA 2, and Falcon). Notably, our method demonstrates particular efficacy in addressing complex speech sentences, acronyms, and domain-specific vocabulary.
Index Terms:
automatic speech recognition, large language models, zero-shot decoding, acoustic modeling, language modeling, MAP inference, iterative decoding.
@@ I Introduction
Automatic Speech Recognition (ASR) has witnessed significant advancements in recent years, driven by the growing capabilities of deep learning and large language models. ASR stands as a cornerstone of intelligent speech technology, enabling machines to understand and transcribe human speech with remarkable precision. It has become an integral part of modern digital experiences, powering applications like virtual assistants, real-time captions on social media platforms, and transcription services for podcasts and meetings.
@@
ASR systems typically comprise three core components: the Acoustic Model (AM), the Language Model (LM), and the decoding mechanism. The AM converts raw audio signals into probabilistic representations of linguistic units, such as phones or sub-word units, capturing the acoustic features necessary for transcription. The LM ensures that the generated sequences are linguistically coherent by predicting the probability of word sequences based on extensive text corpora. The decoding mechanism integrates outputs from both components, utilizing algorithms such as beam search to identify the most likely transcription. Together, these components form the backbone of ASR systems, enabling accurate and contextually relevant transcription of spoken language.
The AM component of ASR systems has undergone significant evolution in recent years. Early approaches were grounded in statistical methods, notably Hidden Markov Models (HMMs) and Gaussian Mixture Models (GMMs) [1, 2]. The emergence of deep learning introduced a major paradigm shift, enabling neural networks to capture complex speech patterns and perform robustly across varying conditions such as noise and speaker accents. A significant departure from traditional HMM-based systems came with the introduction of deep Recurrent Neural Networks (RNNs) [3] and fully convolutional architectures [4], both of which were trained using the Connectionist Temporal Classification (CTC) loss function [5]. In more recent developments, transformer-based models [6], Attention-based Encoder-Decoder (AED) architectures [7], and Recurrent Neural Network Transducers (RNN-T) [8] have become the dominant choices. Additionally, self-supervised learning (SSL) methods such as wav2vec 2.0 [9] and HuBERT [10] have further advanced acoustic modeling by leveraging large-scale unlabeled audio for pretraining, followed by fine-tuning on smaller labeled datasets.
@@
The LM component has traditionally relied on N-gram models, which are probabilistic models predicting the likelihood of a word sequence based on the preceding N-1 words [11]. These models derive their probabilities from extensive text corpora, employing Markov assumptions to simplify calculations by limiting context. Although N-gram models are efficient and interpretable, they face challenges with data sparsity and long-range dependencies, rendering them less effective compared to modern neural language models in ASR.
@@
Recent research has explored methods to enhance traditional N-gram models by utilizing advanced deep learning architectures. One notable approach is the Gated CNN (GCNN) [12], which uses stacked causal convolutions to create a neural language model that effectively captures local context within fixed-length token windows during decoding. Another significant model is the TransformerLM [13], which employs multi-head self-attention to model global dependencies, allowing it to condition on the entire history of hypotheses. These innovations represent a shift towards more sophisticated language modeling techniques that address the limitations of N-gram models.
@@
Large Language Models (LLMs) have achieved remarkable success in various tasks, including machine translation, conversational AI, and text summarization [14, 15]. Pre-trained on diverse textual corpora, LLMs excel at capturing linguistic nuances, contextual relationships, and semantic meanings. As ASR technology approaches human-like accuracy [16], it is only natural to try to replace traditional LMs with LLMs and leverage their advanced contextual understanding to further enhance recognition quality and downstream applications.
@@
The integration of LLMs into ASR systems introduces some challenges [17]. Unlike traditional LMs, LLMs provide a deeper understanding of language and context, enabling ASR systems to handle ambiguities in acoustic signals and domain-specific vocabulary more accurately. Although numerous efforts have been made to incorporate LLMs into ASR systems [17], these approaches typically focus on prompt design, output re-scoring, or architectural modifications. A comprehensive review of these methods is provided in Section II. Notably, however, none of these approaches directly aim to maximize the posterior probability of the word sequence conditioned on the acoustic input. It is worth noting that Spoken Language Models (SLMs) are sometimes implemented with components that resemble both AMs and LLMs. However, their primary goal is not necessarily ASR; instead, they are designed to function as universal speech processing systems, capable of supporting a wide range of downstream tasks [18].
@@
Our goal is to propose a method for integrating the AM and LLM while keeping them separable, allowing each to be independently trained and improved on its own data, and enabling the system to benefit from the strengths of both components without requiring joint optimization.
@@
This work proposes a novel decoding process governed by the LLM, which is formally derived from a decomposition of the Maximum A Posteriori (MAP) estimator of the word sequence given the speech signal. The decoding process is iterative; at each iteration, a set of candidate tokens is sampled from the LLM based on the previously predicted tokens, which are provided as a prompt to the LLM. The candidate tokens are aligned with the speech signal using an AM, and a score is assigned to each alignment. The candidates with the highest combined score from both the AM and the LLM are added to the beam and incorporated into the prompt for the next iteration. The process concludes when all beams reach the end of the sentence or when a predetermined safety horizon is met.
@@
Our method functions in a zero-shot manner, meaning that no additional training is necessary. This approach enables enhancements to the AM or LLM to be implemented independently, directly impacting the ASR output without the need for retraining. However, a notable drawback of our approach is its higher computational demands compared to standard inference, which we will discuss in further detail.
@@
We demonstrate the effectiveness of the proposed method across three speech corpora, each showcasing a distinct speaking style: ALLSTTAR, which features short, simple sentences; WSJ, comprising read speech from newspapers; and TED-LIUM, consisting of presentations by real-world professionals. In the experiments, we used two state-of-the-art acoustic models (wav2vec 2.0, HuBERT) and three LLMs (GPT-2, LLaMA 2, Falcon). The integration of LLM into ASR proves advantageous for more complex speaking styles, enhancing its ability to manage acronyms and specialized vocabulary. Our method shows particular strength in handling complex speech nuances such as acronyms and domain-specific vocabulary
@@ II Related Work
Recent research has explored diverse approaches to integrating LLMs into ASR systems, from post-processing methods to deep architectural changes. These approaches can be grouped into four categories: ASR error correction, N-best list rescoring, prompt engineering, and architecture adaptation. Each represents a trade-off between implementation simplicity and the potential for performance gains, setting the stage for more holistic integration methods.
Prompt Engineering provides a straightforward method for incorporating LLMs into ASR systems. Li et al. [19] explored zero-shot domain adaptation by using domain-specific text prompts with LLMs, integrating second-pass rescoring and LLM fusion with the ASR decoder. Sachdev et al. [20] introduced EvoPrompt, an evolutionary optimization algorithm that systematically refines prompts to improve transcription accuracy with N-best hypotheses. While these methods avoid modifying ASR architectures and require minimal computational resources, they process acoustic and linguistic information independently, limiting joint optimization. Additionally, their effectiveness diminishes in specialized technical domains, motivating further research into integrated solutions.
ASR Error Correction is a post-processing approach that employs LLMs to refine transcription outputs by leveraging their language understanding capabilities, enhancing quality without modifying the ASR system itself [21, 22]. Despite this approach offering simplicity and applicability, it operates solely on the final transcription output, potentially missing valuable acoustic information that could enhance correction accuracy.
N-Best List Rescoring extends error correction by utilizing multiple candidate transcriptions generated by the ASR system. This approach processes N-best lists to identify and address potential errors through multiple output candidates [23, 24]. Recent work has incorporated generative LLMs to map these lists to final transcriptions [25] and evaluated their effectiveness in zero-shot and few-shot settings [26]. Ma et al. [27] further leveraged N-best lists to provide a broader context for error correction, enabling the evaluation of diverse linguistic hypotheses. However, it faces similar limitations, relying on the N-best list’s quality, missing errors outside the candidates, and limiting LLMs’ potential by restricting their use to a post-processing layer instead of integrating them into the ASR system.
Architecture Adaptation involves modifying model architectures to integrate LLMs into ASR systems. Recent methods require varying degrees of architectural changes and retraining. For a comprehensive overview, see the recent survey in [18]; here, we discuss only the most pertinent contributions. Deng et al. [28] proposed Transducer-LLaMA, integrating LLMs into a Factorized Transducer model with vocabulary adaptation, trained using RNN-T and MWER losses. Jia et al. [29] introduced SpeechLLM-XL, a decoder-only model trained on paired audio-text data to predict text tokens auto-regressively, using CTC forced alignment to segment text boundaries. Ma et al. [30] developed SLAM-ASR, incorporating a trainable linear projector between the speech encoder and LLM. Mundnich et al. [31] proposed a lightweight adaptation module for mapping audio representations to LLM token embeddings, facilitating multilingual applications. However, these methods require modifications and retraining, preventing robust, off-the-shelf integration of pre-trained models into ASR systems.
@@ III Background
The speech recognition task involves mapping an acoustic signal to its corresponding textual transcription. Formally, let $\mathbf{X}=(\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{T})$ represent the input audio sequence of length $T$, and $\mathbf{x}_{t}\in\mathbb{R}^{d}$ is a frame of speech, where $t\in[1,T]$ is the frame index. Let ${W}=(w_{1},w_{2},\dots,w_{K})$ denote the corresponding spoken word sequence. We assume that $K$ words were spoken, and $w_{k}\in\mathcal{V}$ for $1\leq k\leq K$ and $\mathcal{V}$ is the vocabulary. The objective is to determine the most probable transcription ${W}^{*}$ given the acoustic input $\mathbf{X}$. In other words, the recognizer should maximize the probability $P({W}|\mathbf{X})$ of the word sequence ${W}$ spoken within $\mathbf{X}$ [32]:
@@
where $P(\mathbf{X}|{W})$ represents the probability of the audio sequence given the word sequence, and it is referred to as the acoustic model, and the probability $P({W})$ represents the probability of the word sequence and is called language model. The process of identifying the word sequence ${W}^{*}$ that maximizes the MAP estimate is known as decoding. We provide a brief overview of each component, focusing on aspects pertinent to our proposed approach.
@@ III-A Acoustic model
Traditionally, the acoustic model $P(\mathbf{X}|{W})$ was estimated using HMMs. In its standard form, an individual HMM is defined for each phone, and word-level HMMs are constructed by sequentially concatenating the phone-specific models. To represent an entire utterance, the HMMs corresponding to each word in the sequence ${W}$ are further concatenated, yielding a single HMM that models the full spoken input [33]. The model presumes that the probability of transitioning to a particular state depends solely on the current state, independent of the previous states. Additionally, the output probability of a speech frame is considered conditionally independent of prior or future frames, given the current state [32]. HMM-based acoustic models lag behind modern transformer-based models.
Our work focuses on modern recent acoustic models, based on self-supervised learning representations, such as wav2vec 2.0 [9] and HuBERT [10]. These models are initially trained on raw, unlabeled audio to learn contextualized representations of speech. Then, they are fine-tuned on transcribed speech data using the Connectionist Temporal Classification (CTC) loss function [5], enabling them to serve as effective acoustic models for automatic speech recognition. This loss enables them to operate directly on character sequences, rather than phonetic states, without requiring explicit alignment between the input audio and target output. As a result, these models adopt a fundamentally different formulation of the acoustic modeling task. Instead of estimating $P(\mathbf{X}|{W})$, they estimate the posterior probabilities $P({Y}|\mathbf{X})$, where ${Y}$ represents a sequence of $U$ characters, including the blank token $\epsilon$. Specifically, ${Y}=(y_{1},\ldots,y_{U})$, with $y_{i}\in\{a,\ldots,z,\epsilon\}$ [9, 10].
We note that the primary focus of this work is to propose a mechanism for integrating an LLM as the language model component alongside a given acoustic model. As such, we do not address fully end-to-end supervised transformer-based ASR systems, such as Whisper [34], which do not readily separate into distinct acoustic and language modeling components. We leave the exploration of such models for future work.
@@ III-B Language model
The language model (LM) plays a critical role in ensuring that the recognized text is linguistically plausible and contextually coherent. Its primary goal is to assign a probability to a given sequence of words and guide the decoding process toward the most likely interpretation of the spoken input. The language model probability $P({W})$ is often decomposed as
@@
where $w_{1}^{n-1}=(w_{1},\ldots,w_{n-1})$. The most commonly used language model in ASR is the N-gram language model [10, 9]. This probabilistic model estimates the likelihood of a word sequence by assuming that each word depends only on the preceding N-1 words [11]. A key limitation of N-gram models is their inability to capture long-range dependencies due to their fixed-size context window. More recent approaches employ advanced neural language models, such as Gated Convolutional Neural Networks (GCNNs) [12], a non-recurrent architecture that uses stacked convolutions and a simplified gating mechanism to efficiently model long-range dependencies, and TransformerLM [13], which is based on a decoder-only transformer architecture.
LLMs, such as GPT-2 [14] and LLaMA 2 [15], are deep neural networks based on the transformer architecture, trained to predict the next token (typically a sub-word unit) given a sequence of preceding tokens. Although LLMs follow the same conditional formulation as in (4), they differ fundamentally from N-gram models. LLMs are trained using the cross-entropy loss over massive text corpora to learn the parameters of a transformer-based architecture. This training often results in a high norm of the output logit weights, which causes the softmax function to produce an overly sharp, overconfident probability distribution [35].
The effective use of LLMs as a language model for ASR is still an open problem. While LLM models predict the probability of the next token given the sequence of previous tokens, the resulting probability is not calibrated due to the way they are trained [36]. Furthermore, ASR systems require language models that support incremental scoring, meaning they can evaluate partial word sequences (prefixes), manage multiple parallel hypotheses during beam search, and synchronize closely with the acoustic model’s step-by-step output [32]. However, LLMs are built to process complete sequences in a forward-only manner, making them ill-suited for this role. They lack native support for prefix scoring, cannot easily backtrack or compare multiple evolving hypotheses, and perform best when given the entire context, which is often unavailable in real-time ASR decoding. Since LLMs cannot be directly integrated into the acoustic model, we propose to incorporate them through the decoding process.

### v2

# Large Language Model Guided Decoding for Self-Supervised Speech Recognition

@@ Abstract
Self-supervised automatic speech recognition (SSL-ASR) is an ASR approach that uses speech encoders pretrained on large amounts of unlabeled audio (e.g., wav2vec2.0 or HuBERT) and then fine-tunes them with limited labeled data to perform transcription. Decoding is usually performed with a CTC decoder, whose hypotheses are scored and refined using an external language model (LM), typically an n-gram or neural LM, which guides beam search to produce the final transcription. Using Large Language Models (LLMs) as external LMs remains a challenge, as their word probabilities are overly confident.
The proposed method integrates an LLM with an SSL acoustic model by using the LLM’s decoding mechanism to generate a set of candidate next tokens. For each candidate, the SSL model provides an acoustic score by aligning it to the input acoustics of the SSL model. A combined acoustic and LLM score is then calculated based on decomposing the MAP estimator of words given the acoustic signal. The tokens with the highest combined scores are maintained in a beam, which is then used to proceed to the next decoding step. We illustrate the effectiveness of our method through a comprehensive comparison with the current state-of-the-art LLM-based decoding, post-processing, and error-correcting methods across multiple datasets. Our approach proves particularly effective when processing challenging inputs such as complex speech sentences, acronyms, and domain-specific vocabulary.
@@ I Introduction
Self-supervised automatic speech recognition (SSL-ASR) refers to ASR systems that use speech encoders pretrained on unlabeled audio using self-supervised learning methods such as wav2vec 2.0 [baevski2020wav2vec], HuBERT [hsu2021hubert], or WavLM [chen2022wavlm]. These pretrained encoders are then fine-tuned with a small amount of labeled data—typically using CTC or a light decoder—to perform speech recognition. SSL-ASR is especially effective in low-resource or domain-adaptation settings because it decouples representation learning from the need for large supervised datasets. Although supervised and weakly supervised approaches such as Whisper [radford2023whisper] have been shown to approach human-level transcription accuracy [kim2024automatic], there remain scenarios, particularly in low-resource conditions, domain adaptation, or child speech, where SSL-based ASR models are more suitable and often preferred [barcovschi2023comparative, anidjar2024whisper, rouditchenko2023comparison, adnan2025one].
@@
SSL-ASR systems generally consist of three core components: the Acoustic Model (AM), the Language Model (LM), and the decoding mechanism. The AM converts raw audio signals into probabilistic representations of linguistic units, such as phones or sub-word units, capturing the acoustic features necessary for transcription. In this paper, we restrict ourselves to SSL-based models. The LM component ensures that the generated sequences are linguistically coherent by predicting the probability of word sequences based on extensive text corpora. The decoding mechanism integrates outputs from both components, utilizing algorithms such as beam search to identify the most likely transcription.
@@
The LM component has traditionally relied on N-gram models, which are probabilistic models predicting the likelihood of a word sequence based on the preceding N-1 words [jm3]. These models derive their probabilities from extensive text corpora, employing Markov assumptions to simplify calculations by limiting context. Although N-gram models are efficient and interpretable, they face challenges with data sparsity and long-range dependencies, rendering them less effective compared to modern neural language models in ASR.
@@
Advances beyond traditional N-gram models have led to substantial improvements in language modeling, driven largely by the adoption of more expressive deep learning architectures. One notable approach is the Gated CNN (GCNN) [dauphin2017language], which uses stacked causal convolutions to create a neural language model that effectively captures local context within fixed-length token windows during decoding. Another significant model is the TransformerLM [baevski2018adaptive], which employs multi-head self-attention to model global dependencies, allowing it to condition on the entire history of hypotheses. These developments mark a clear transition toward more powerful and flexible language modeling paradigms that overcome key limitations of conventional N-gram approaches.
@@
Large Language Models (LLMs) have achieved remarkable success in various tasks, including machine translation, conversational AI, and text summarization [radford2019language, touvron2023llama]. Pre-trained on diverse textual corpora, LLMs excel at capturing linguistic nuances, contextual relationships, and semantic meanings. It is only natural to try to replace traditional LMs with LLMs and leverage their advanced contextual understanding to further enhance recognition quality and downstream applications.
@@
The integration of LLMs into SSL-ASR systems introduces some challenges [toshniwal2018comparison]. Unlike traditional LMs, LLMs provide a deeper understanding of language and context, and can enable SSL-ASR systems to handle ambiguities in acoustic signals and domain-specific vocabulary more accurately. Although numerous efforts have been made to incorporate LLMs into SSL-ASR systems [toshniwal2018comparison, ma2023can, tur2024progres], these approaches typically focus on prompt design, output re-scoring, error-correction, or architectural modifications. A comprehensive review of these methods is provided in Section II. Notably, however, none of these approaches integrates the LLM with the AM in a manner that enables mathematically principled inference through direct maximization of the posterior probability (MAP) of the word sequence given the acoustic input. It is worth noting that Spoken Language Models (SLMs) are constructed with components that are functionally analogous to both AMs and LLMs. However, their primary goal is not necessarily ASR; instead, they are designed to function as universal speech processing systems, capable of supporting a wide range of downstream tasks [arora2025landscape].
@@
Our goal is to propose a method for integrating LLMs into SSL-based acoustic models while keeping them separable, allowing each to be independently trained and improved on its own data, and enabling the system to benefit from the strengths of both components without requiring joint optimization.
@@
We propose a new decoding mechanism guided by the LLM, formally derived from a decomposition of the MAP estimator for the word sequence given the speech signal. The decoding process is iterative. In each iteration, the LLM samples a set of candidate tokens conditioned on the previously predicted tokens. The candidate tokens are aligned with the speech signal using the SSL model, which outputs an acoustic score for each alignment. The candidates with the highest combined score from both the SSL model and the LLM are added to the beam and incorporated into the prompt for the next iteration. The process concludes when all beams reach the end of the sentence or when a predetermined safety horizon is met.
@@
Our method operates in a zero-shot manner, meaning no additional training is required. This approach enables enhancements to the SSL model or LLM to be implemented independently, directly affecting ASR output without retraining. However, a notable drawback of our approach is its higher computational demands compared to standard inference, which we will discuss in further detail.
@@
We demonstrate the effectiveness of the proposed method across three speech corpora, each showcasing a distinct speaking style: ALLSTTAR, which features short, simple sentences; WSJ, comprising read speech from newspapers; and TED-LIUM, consisting of presentations by real-world professionals. In the experiments, we used two state-of-the-art acoustic models (wav2vec 2.0, HuBERT) and three LLMs (GPT-2, LLaMA 2, Falcon). The integration of LLM into ASR proves advantageous for more complex speaking styles, enhancing its ability to manage acronyms and specialized vocabulary. Lastly, we compare out method with other decoding techniques, post-processing methods and LLM-based error-correcting methods. Our method shows particular strength in handling complex speech nuances such as acronyms and domain-specific vocabulary
@@ II Related Work
In traditional ASR systems, decoding is commonly formulated as a MAP search that integrates acoustic and language model scores. In the context of SSL-CTC acoustic models, this is most commonly realized through shallow fusion with external language models, including traditional N-gram models [jm3] as well as neural approaches such as Gated CNNs [dauphin2017language] and Transformer-based language models [baevski2018adaptive].
The main line of work applies LLMs in a post-hoc manner, operating on the final ASR transcription to refine and improve the output. One such approach is N-best rescoring, where a language model assigns scores to multiple candidate hypotheses produced by the ASR system and selects the most probable one [chelba2012large, xu2018pruned, udagawa2022effect, xu2022rescorebert]. Another post-hoc approach is error correction, in which language models directly correct or refine the ASR output. [ma2023can] showed that general-purpose LLMs can perform ASR error correction in a zero-shot or one-shot setting by prompting the model with the ASR N-best hypotheses. They also demonstrated a supervised variant based on a fine-tuned T5 model operating on the same N-best input. [ma2024asr] extended this work with additional post-hoc correction strategies, including constrained variants that operate over N-best lists or lattices. [tur2024progres] proposed a hybrid post-hoc approach that augments the ASR N-best list with LLM-generated hypotheses and then jointly rescores both original and generated candidates using additional ASR confidence scores and LLM sequence scoring.
A complementary approach integrates LLMs into ASR systems by modifying the underlying architecture or through joint training of speech-text models. Some approaches introduce trainable projection layers that connect SSL encoder representations to an LLM [ma2024embarrassingly, mundnich2024zero]. Other methods train multimodal speech-language models used for multiple downstream tasks [deng2024transducer, jia2024efficient]. Although effective, these systems require supervised training or architectural changes and are therefore out of scope for this work.
@@ III Background
The speech recognition task involves mapping an acoustic signal to its corresponding textual transcription. Formally, let $\mathbf{X}=(\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{T})$ represent the input audio sequence of length $T$, and $\mathbf{x}_{t}\in\mathbb{R}^{d}$ is a frame of speech, where $t\in[1,T]$ is the frame index. Let ${W}=(w_{1},w_{2},\dots,w_{K})$ denote the corresponding spoken word sequence. We assume that $K$ words were spoken, and $w_{k}\in\mathcal{V}$ for $1\leq k\leq K$ and $\mathcal{V}$ is the vocabulary. The objective is to determine the most probable transcription ${W}^{*}$ given the acoustic input $\mathbf{X}$. In other words, the recognizer should maximize the probability $P({W}|\mathbf{X})$ of the word sequence W{W} spoken within $\mathbf{X}$ [Jelinek1998]:
@@
where $P(\mathbf{X}|{W})$ represents the probability of the audio sequence given the word sequence, and it is referred to as the acoustic model, and the probability P(W)P({W}) represents the probability of the word sequence and is referred to as the language model. The process of identifying the word sequence ${W}^{*}$ that maximizes the MAP estimate is known as decoding. We provide a brief overview of each component, focusing on aspects pertinent to our proposed approach.
@@ III-A Acoustic model
Traditionally, the acoustic model $P(\mathbf{X}|{W})$ was estimated using HMMs. In its standard form, an individual HMM is defined for each phone, and word-level HMMs are constructed by sequentially concatenating the phone-specific models. To represent an entire utterance, the HMMs corresponding to each word in the sequence W{W} are further concatenated, yielding a single HMM that models the full spoken input [rabiner1993fundamentals]. The model presumes that the probability of transitioning to a particular state depends solely on the current state, independent of the previous states. Additionally, the output probability of a speech frame is considered conditionally independent of prior or future frames, given the current state [Jelinek1998]. HMM-based acoustic models lag behind modern transformer-based models.
Our work focuses on acoustic modeling approaches based on self-supervised representations, including wav2vec 2.0 [baevski2020wav2vec] and HuBERT [hsu2021hubert]. These models are initially trained on raw, unlabeled audio to learn contextualized representations of speech. Then, they are fine-tuned on transcribed speech data using the Connectionist Temporal Classification (CTC) loss function [graves2006connectionist], enabling them to serve as effective acoustic models for automatic speech recognition. This loss enables them to operate directly on character sequences, rather than phonetic states, without requiring explicit alignment between the input audio and target output. As a result, these models adopt a fundamentally different formulation of the acoustic modeling task. Instead of estimating $P(\mathbf{X}|{W})$, they estimate the posterior probabilities $P({Y}|\mathbf{X})$, where Y{Y} represents a sequence of $U$ characters, including the blank token $\epsilon$. Specifically, ${Y}=(y_{1},\ldots,y_{U})$, with $y_{i}\in\{a,\ldots,z,\epsilon\}$ [baevski2020wav2vec, hsu2021hubert].
@@ III-B Language model
The language model plays a critical role in ensuring that the recognized text is linguistically plausible and contextually coherent. Its primary goal is to assign a probability to a given sequence of words and guide the decoding process toward the most likely interpretation of the spoken input. The language model probability P(W)P({W}) is often decomposed as
@@
where $w_{1}^{n-1}=(w_{1},\ldots,w_{n-1})$. The most commonly used language model used with SSL acoustic models is the N-gram language model [baevski2020wav2vec, hsu2021hubert], but GCNNs [dauphin2017language], and TransformerLM [baevski2018adaptive] are used.
LLMs, such as GPT-2 [radford2019language] and LLaMA 2 [touvron2023llama], are deep neural networks based on the transformer architecture, trained to predict the next token (typically a sub-word unit) given a sequence of preceding tokens. Although LLMs follow the same conditional formulation as in (4), they differ fundamentally from N-gram models. LLMs are trained using the cross-entropy loss over massive text corpora to learn the parameters of a transformer-based architecture. This training often results in a high norm of the output logit weights, which causes the softmax function to produce an overly sharp, overconfident probability distribution [pmlr-v162-wei22d].
Despite their strong generative capabilities, the effective use of LLMs as language models for SSL acoustic models remains challenging. While LLM models predict the probability of the next token given the sequence of previous tokens, the resulting probability is not calibrated due to the way they are trained [Bridle1989]. Furthermore, ASR systems require language models that support incremental scoring, meaning they can evaluate partial word sequences (prefixes), manage multiple parallel hypotheses during beam search, and synchronize closely with the acoustic model’s step-by-step output [Jelinek1998]. However, LLMs are built to process complete sequences in a forward-only manner, making them ill-suited for this role. They lack native support for prefix scoring, cannot easily backtrack or compare multiple evolving hypotheses, and perform best when given the entire context, which is often unavailable in real-time ASR decoding. Since LLMs cannot be directly integrated into the acoustic model, we propose to incorporate them through the decoding process.